# Understanding Before Reasoning: Enhancing Chain-of-Thought with Iterative Summarization Pre-Prompting

### v1

@@ 3.2 Reliability Rating of Information Pairs
  • Opinion-Based Judgment Prompt: Prompts the LLM to generate opinion-based judgments (e.g., absolutely reliable, moderately reliable, weakly reliable, unreliable), which can be converted into numerical values $v$ (1, 0.67, 0.33, 0).
To construct the prompt $T$, we provide stepwise evaluation examples (similar to question answering with rationales) for each instance. Reliability rating prompt takes different forms depending on the specific problem, enabling the LLM to assess the information pair and assign an appropriate value based on its judgment.
3.3 Iterative Summarization for Knowledge Understanding
@@
At this step, the goal is to summarize the information pairs from adaptive extraction, refining the final knowledge that can aid in solving the problem. Among the given $n$ information pairs, we select the two with the lowest reliability scores for merging, rather than those with higher scores. The transition from low to high reliability scores reflects the evolution of knowledge from incomplete to complete information, representing a process of deepening understanding. Adaptive extraction focuses on identifying explicit information, while summarization goes beyond this by inferring implicit knowledge from the previously explicit information, thereby providing deeper insights. In this way, the selection of lower-scoring information pairs allows for the continuous refinement and expansion of knowledge. The iterative process facilitates a more comprehensive understanding of the problem, ensuring that both explicit and implicit knowledge are effectively utilized. As illustrated in Figure 2, specifically, we provide the input context $x$ and the two low-scored information pairs $IP_{a}$ and $IP_{b}$, and reason through the LLM to generate a new information pair $IP_{\text{new}}$, represented as:
$IP_{\text{new}}=\arg\max_{\begin{subarray}{c}IP_{\text{new}}\subseteq IP_{a}\cup IP_{b}\end{subarray}}\prod_{i=1}^{k}p_{i}(IP_{\text{new}}|x),$ (2)
where $\prod_{i=1}^{k}p_{i}$ represents the product of probabilities $P_{i}$ corresponding to all elements in $IP_{\text{new}}$. The new information pair $IP_{\text{new}}$ replaces the original two and is further evaluated for reliability. The process of summarization and scoring continues through multiple iterations until the LLM generates the final information pair, and then the iterative process ends. For example, as shown in Figure 3, four information pairs related to the problem have been obtained, and through reliability scoring, "Bolts" and "Robe" are found to have the lowest scores. Then, by using designed prompts to guide the LLM, a new information pair "Fiber requirements for robe" is generated, with knowledge descriptions: "1. A robe requires 2 bolts of blue fiber and half that amount of white fiber." and "2. The total bolts needed for the robe can be calculated by adding the bolts of blue and white fiber." This new information will be added to the list of information pairs, and "Bolts" and "Robe" will be removed. The information pairs in the list are then re-evaluated to facilitate finding new low-scored information pairs in the next iteration. The new prompt is then passed to the LLM, and iterations continue until the model returns a unique information pair.
3.4 Question Answering
@@
We treat the reasoning process of LLMs as an autoregressive generation task. Typically, the input context $x$ consists of two parts: the question $Q$ and the prompt $T$. Given $x$, the model needs to generate the final result $y$. To generate a reasonable $y$, the LLM needs to leverage CoT methods and reason correctly through $z$ as an intermediate step. We define the predictive probability formula as follows:
@@
where the LLM can be used to compute each condition $p(y\mid x)$ by incorporating the conditional variables ($Q$ and $T$) as part of its input.
Building on the foundation, $\text{ISP}^{2}$ enhances reasoning capabilities by integrating information extraction with cognitive summarization. Compared to standard prompting methods, $\text{ISP}^{2}$ is a pre-prompting method that focuses on understanding the problem. It employs an iterative summarization process before inference, progressively refining information pairs based on the given question. $\text{ISP}^{2}$ not only extracts relevant information but also deeply comprehends the context and nuances of the problem at hand. The method allows the LLM to generate more accurate and contextually relevant information pairs for the final reasoning step. Ultimately, the unique information pair $IP$ is combined with the question for the final reasoning process. Formally, it can be represented as:
@@
where $IP$ represents the unique information pair generated through iterative refinement. We present examples and the workflow in Figures 2 and 3, respectively, and provide the complete generation results in the appendix. To further improve the quality and accuracy of results in complex problem scenarios, $\text{ISP}^{2}$ uses specialized prompts such as $\text{ISP}^{2}$-CoT and $\text{ISP}^{2}$-Complex CoT prompts for generating final answers.
4 Experiments
@@ 4.1 Experimental Setup
We evaluate $\text{ISP}^{2}$ on six datasets with diverse input formats. Extensive experiments are conducted across these datasets to demonstrate the universality of $\text{ISP}^{2}$ prompts. The Table 1 provides relevant information about the datasets used in our experiment, detailing the data source, task type, answer type, number of prompt samples, and total test samples for each dataset.
  • Arithmetic Reasoning: The following four arithmetic reasoning benchmarks are widely recognized and considered in the field: AddSub Hosseini et al. (2014), which includes math word problems on addition and subtraction tailored for third to fifth graders; SVAMP Patel, Bhattamishra, and Goyal (2021), known for its math word problems with diverse structures; AQuA Ling et al. (2017), which focuses on algebraic word problems; GSM8K Cobbe et al. (2021), a published benchmark that features grade-school math problems.
  • Commonsense Reasoning: StrategyQA (SQA; Geva et al., 2021) and CommonsenseQA (CSQA; Talmor et al., 2019) are utilized for commonsense tasks. CSQA comprises questions that require a variety of commonsense knowledge, while StrategyQA includes questions that necessitate multi-step reasoning.
@@
GSM8K Cobbe et al. (2021) Arithmetic Number 4 1,319
AddSub Hosseini et al. (2014) Arithmetic Number 4 395
SVAMP Patel, Bhattamishra, and Goyal (2021) Arithmetic Number 4 1,000
AQuA Ling et al. (2017) Arithmetic Multi-choice 4 254
StrategyQA Geva et al. (2021) Commonsense True/False 4 2,290
CommonsenseQA Talmor et al. (2019) Commonsense Multi-choice 4 1,221
Table 1: Overview of datasets utilized in experiments, where “Number” represents the number of sampled datasets, and “Example” is the number of prompt examples in the same dataset.
@@
To effectively evaluate our method, we assess $\text{ISP}^{2}$ performance on three baseline prompting methods: CoT Wei et al. (2022), Complex CoT Fu et al. (2023), and Self Consistency Wang et al. (2023). CoT predicts answers by generating explanations and steps, allowing the model to solve problems through explicit reasoning processes, which makes the decision-making more transparent. Complex CoT utilizes a complexity-based strategy, which breaks down complex problems into smaller parts, thereby improving reliability in complex scenarios. Self Consistency generates multiple chains of thought and selects the highest-voted result as the final outcome through voting. Additionally, all experiments are conducted in a few-shot setting without training or fine-tuning the LLMs.
@@
We primarily use GPT-3.5111The GPT-3.5 API model ”gpt-3.5-turbo-0125” is uniformly used in this paper, and it is available at: https://platform.openai.com/docs/models#gpt-3-5-turbo and LLaMA2-7B, LLaMA2-13B [^2] LLaMA2-7B and LLaMA2-13B, developed by Meta, have 7 billion and 13 billion parameters, respectively. The hf versions are available at: https://huggingface.co/meta-llama/Llama-2-7b-hf and https://huggingface.co/meta-llama/Llama-2-13b-hf. for testing, and our decoding strategy employs greedy decoding with the temperature set to 0, thus producing deterministic outputs. For the few-shot setting, we commonly use four samples as exemplars, depending on the difficulty of the dataset. We integrate $\text{ISP}^{2}$ into these baseline methods to evaluate its impact, denoted as $\text{ISP}^{2}$-CoT, $\text{ISP}^{2}$-CoT@5, $\text{ISP}^{2}$-ComCoT, and $\text{ISP}^{2}$-ComCoT@5. Here, $\text{ISP}^{2}$-CoT and $\text{ISP}^{2}$-ComCoT represent $\text{ISP}^{2}$ combined with CoT and Complex CoT, respectively. The "@5" notation indicates the use of Self Consistency by retrieving five reasoning chains to perform majority voting, enhancing the robustness of the final output. Additionally, for different tasks, we design answer format instructions in the prompts to regulate the structure of the final answer, facilitating precise answer extraction.
4.2 Main Results
The main results are presented in Table 2, Table 3, and Table 4. Compared with the SOTA method that also functions as a plug-in, $\text{ISP}^{2}$ demonstrates considerable advantages. Significant improvements have been achieved on benchmarks such for GPT-3.5, LLaMA2-13B, and LLaMA2-7B. The average improvement rates were 7.1% for GPT-3.5, 8.1% for LLaMA2-13B, and 12.4% for LLaMA2-7B. These results indicate that by processing $\text{ISP}^{2}$, LLM can better understand the essence of the problems and enhance its performance.
Model Method Dataset Average
AddSub SVAMP GSM8K AQuA CSQA SQA
GPT-3.5 Turbo CoT 81.2 81.2 76.2 58.7 75.5 69.7 73.75
@@ 4.2 Main Results
+7.4 +7.5 +7.2 +5.7 +3.2 +1.7 +5.68
CoT@5 82.3 85.2 80.8 66.5 76.7 71.2 77.12
$\text{ISP}^{2}$-CoT@5 93.9 90.1 84.8 72.7 81.0 73.2 82.62
+11.6 +4.9 +4.0 +6.2 +4.3 +2.0 +5.50
LLaMA2-13B CoT 38.1 36.3 16.7 16.6 46.9 61.3 35.98
$\text{ISP}^{2}$-CoT 60.0 39.2 19.3 20.4 49.2 61.4 41.58
+21.9 +2.9 +2.6 +3.8 +2.3 +0.1 +5.60
CoT@5 47.8 47.8 17.8 24.4 53.9 61.4 42.18
$\text{ISP}^{2}$-CoT@5 64.6 49.7 21.9 25.4 53.7 64.5 46.63
+16.8 +1.9 +4.1 +1.0 -0.2 +3.1 +4.45
LLaMA2-7B CoT 29.4 30.7 7.4 19.7 27.2 54.7 28.18
$\text{ISP}^{2}$-CoT 35.2 39.1 8.3 21.6 28.4 56.8 31.57
+5.8 +8.4 +0.9 +1.9 +1.2 +2.1 +3.39
CoT@5 41.8 33.4 7.8 21.6 29.1 57.4 31.85
$\text{ISP}^{2}$-CoT@5 43.9 42.5 11.4 29.9 30.1 58.5 36.05
+2.1 +9.1 +3.6 +8.3 +1.0 +0.9 +4.20
Table 2: $\text{ISP}^{2}$ can help to improve the performance when applied to different LLMs and prompting methods. @5 utilizes Self Consistency by retrieving five CoT reasoning chains to make majority votes.
@@
GPT-3.5 Turbo ComCoT 82.7 80.1 79.3 57.8 74.98
@@
LLaMA2-13B ComCoT 32.2 33.9 15.5 19.7 25.33
@@
Table 3: $\text{ISP}^{2}$ can help to improve the performance when applied to different LLMs and prompting methods. @5 utilizes Self Consistency by retrieving five Complex CoT reasoning chains to make majority votes. "ComCoT" stands for Complex CoT.
@@
Table 2 and Table 3 reports performance on the Math Word Problem (MWP) task, where our method achieves significant performance improvements across various mathematical subdomains, surpassing ComplexCoT by $5.4\%$ on GPT 3.5. Notably, the enhancement is particularly significant when combined with Self Consistency, and the robustness and accuracy of $\text{ISP}^{2}$ on different models are evident. The AddSub dataset focuses on basic mathematical operations, and for models like GPT and Llama, they already possess sufficient capabilities to solve most problems. However, problems that are not successfully solved usually contain misleading information, and $\text{ISP}^{2}$ can avoid the influence of misleading information by filtering out unnecessary data in adaptive extraction, thus enabling more successful problem solving on LLMs. SVAMP and GSM8K are more advanced MWP datasets that focus on generalization capabilities in arithmetic and more complex algebra and geometry problems. The Base Prompt method shows a notable decrease compared to AddSub on two datasets. Interestingly, we observe that the improvements of $\text{ISP}^{2}$ on SVAMP and GSM8K are not less than those on AddSub. $\text{ISP}^{2}$ can correct wrong directions through adaptive extraction and prevent misleading information from contaminating the summary process. Unlike the first three datasets, AQuA is an algebraic multiple-choice dataset, which not only requires the model to solve the problem but also to select the correct answer from multiple options. The format demands higher judgment capabilities from the model. $\text{ISP}^{2}$ facilitates further computations by allowing algebraic information pairs to be stored in formulaic forms. $\text{ISP}^{2}$ demonstrates good performance on mathematical datasets, enhancing the effectiveness of solutions by revealing key data essential for problem resolution, thus improving the performance of both ComplexCoT and Self Consistency.
@@
Notably, as detailed in Table 4, compared to the current SOTA plug-and-play methods in mathematics, our approach performs well, reaching new best levels in AddSub and SVAMP, and second-best levels in GSM8K and AQuA. The main reason is that $\text{ISP}^{2}$ continuously improves solutions by selectively gathering previous descriptions of problem space, allowing it to accurately resolve issues. Instead of relying on multiple rounds of interactive reasoning, repetitive reading, or relationship extraction to obtain crucial information, solving many complex mathematical problems emphasizes the thought process and information filtering. By focusing on these aspects, misleading information is excluded, and effective problem-solving strategies are developed. It enables $\text{ISP}^{2}$ to achieve significant improvements in mathematical reasoning.
GPT-3.5 turbo Prompt AddSub SVAMP GSM8K AQuA CSQA SQA Average
CoT Wei et al. (2022) 81.2 81.2 76.2 58.7 75.5 69.7 73.75
PHP-CoT Zheng et al. (2024) 86.1 83.1 84.6 65.4 76.2 69.2 77.43
RE2-CoT Xu et al. (2024) 82.7 84.9 81.2 63.3 77.9 67.1 76.60
ERA-CoT Liu et al. (2024) 83.8 82.2 80.2 56.9 83.2 71.4 76.28
$\text{ISP}^{2}$-CoT 88.6 88.7 83.4 64.4 78.7 71.4 79.43
Table 4: The comparison results of existing SOTA plug-and-play methods on GPT-3.5. We use accuracy as the evaluation metric.The best result is highlighted in bold, and the second best is underlined.
@@
As shown in Table 2, for the StrategyQA and CommonsenseQA datasets, $\text{ISP}^{2}$ has increased performance by $4.9\%$ and $2.8\%$, respectively. Additionally, Self Consistency maintains an orthogonal relationship with $\text{ISP}^{2}$, both enhancing the performance of CoT, enabling it to better understand common sense and make more accurate choices in commonsense reasoning. There is a significant performance improvement in the CSQA dataset, where many hidden pieces of information exist within the problems, and relying solely on explicit knowledge is insufficient for accurate responses. Notably, in StrategyQA with the LLaMA2 model, $\text{ISP}^{2}$ did not demonstrate significant improvement. Smaller parameter LLMs have a tendency to include irrelevant text when forming information pairs. Consequently, this inclusion leads to the formation of continuous erroneous reasoning chains. Such errors can impede the ability of $\text{ISP}^{2}$ to effectively enhance performance. However, most experiments still show that $\text{ISP}^{2}$ outperforms most base prompts in commonsense datasets, demonstrating strong capability of $\text{ISP}^{2}$ to guide CoT. When compared with existing SOTA methods, it performs comparably to ERA in SQA, which excels at controlling the reasoning process using entity relations, and also maintains second-best performance in CSQA.
4.3 Ablation Studies

### v2

@@ III-B Reliability Rating of Information Pairs
  • Opinion-Based Judgement Prompt: Prompts the LLM to generate opinion-based judgments (e.g., absolutely reliable, moderately reliable, weakly reliable, unreliable), which can be converted into numerical values $v$ (1, 0.67, 0.33, 0).
To construct the prompt $T$, we provide stepwise evaluation examples (similar to question answering with rationales) for each instance. Reliability rating prompt takes different forms depending on the specific problem, enabling the LLM to assess the information pair and assign an appropriate value based on its judgment. In mathematical tasks, the reliability rating places greater emphasis on logical reasoning and the correctness of derivations, as these are critical for ensuring accurate solutions. In contrast, for commonsense reasoning or general knowledge tasks, the focus shifts toward coherence, relevance, and the ability to provide intuitive and contextually appropriate responses.
III-C Iterative Summarization for Knowledge Understanding
@@
To transform fragmented information into deeply integrated knowledge, we propose an iterative approach that adaptively summarizes and refines the extracted information pairs, progressively distilling knowledge that ultimately contributes to effective problem-solving. It operates through three recurrent operations: selection of two information pairs with minimal reliability scores, generative synthesis of composite knowledge, and reliability rating of the new information pair.
Figure 2 illustrates the computational pipeline, while Figure 3 demonstrates its application in a mathematical dataset. The core of the iterative process lies in summarizing through the merging of two low-reliability information pairs. Among the given $n$ information pairs, the two with the lowest reliability scores are selected for merging in each iteration, rather than those with higher scores. The transition from low to high reliability scores reflects the evolution of knowledge from incomplete to complete information, representing a process of deepening understanding. For example, in the initial stage, the algorithm may select the information pairs ”Bolts” (score: 0.42) and ”Robe” (score: 0.37). Unlike traditional methods that favor high scores inputs, this strategy leverages the insight that low-reliability pairs often contain complementary fragments of knowledge. Merging such pairs can activate latent relationships and enable implicit inference. Combining the pairs <Bolts, “1. Units of measurement for the fiber required. 2. The quantity needed to create the robe” > and <Robe, “1. Requires a specific amount of blue and white fiber” >, a tailored prompt guides the LLM to generate a new information pair: <Total bolts, “1. A robe requires 2 bolts of blue fiber and 1 bolt of white fiber. 2. The total bolts needed for the robe is the sum of blue and white fiber bolts” >. The reliability score of the new pair increases to 0.78, indicating the successful transformation from explicit data to refined implicit knowledge. The merging operation follows a probabilistic maximization principle, where the new information pair $IP_{\text{new}}$ is defined as:
$IP_{\text{new}}=\arg\max_{IP\subseteq IP_{a}\cup IP_{b}}\prod_{k=1}^{m}p_{k}(IP\mid x)$ (3)
where $p_{k}$ denotes the contextual probability of the $k$-th knowledge element given the context $x$. After the merge, the new information pair replaces the original two pairs $IP_{a}$ and $IP_{b}$, and the reliability scores of the entire list are recalculated. The process continues iteratively until it meets the following termination criteria: the list converges to a single information pair, regardless of its reliability score. It ensures termination either when no further merging is possible.
III-D Question Answering
ISP2-CoT Reasoning Problem: A person travels at 20 km/h and reaches the destination in 2.5 hours. Find the distance. Options: A) 53 km, B) 55 km, C) 52 km, D) 60 km, E) 50 km Information Pair
Velocity:
1. The person is traveling at a constant speed of 20 km/hr.
2. The travel time to reach the destination is 2.5 hours. 3. The distance traveled can be found by multiplying the speed by the travel time. Let’s think step by step.
The distance that the person traveled would have been 20 km/hr * 2.5 hrs = 50 km.
Answer: E (50 km)
@@ III-D Question Answering
We treat the reasoning process of LLMs as an autoregressive generation task. Typically, the input context $x$ consists of two parts: the question $Q$ and the prompt $T$. Given $x$, the model needs to generate the final result $y$. To generate a reasonable $y$, the LLM needs to leverage CoT and reason correctly through $z$ as an intermediate step. We define the predictive probability formula as follows:
@@
where $p(y\mid x,z)$ represents the probability of generating the final result $y$ given both the input context $x$ and the intermediate reasoning step $z$, and $p(z\mid x)$ represents the probability of generating the intermediate reasoning step $z$ given the input context $x$. $y$ and $z$ are conditionally independent given $x$. The intermediate reasoning step $z$ can be viewed as independently generated from the final result $y$, conditioned on the input $x$. In practice, this conditional independence reflects the modular nature of the reasoning process, where each step contributes to the overall generation without direct dependence on subsequent steps.
Building on the foundation, ISP2 enhances reasoning capabilities by integrating information extraction with cognitive summarization. It employs an iterative summarization process before inference, progressively refining information pairs based on the given question. ISP2 not only extracts relevant information but also deeply comprehends the context and nuances of the problem at hand. The method allows the LLM to generate more accurate and contextually relevant information pairs for the final reasoning step. Ultimately, the unique information pair IP is combined with the question for the final reasoning process. Formally, it can be represented as:
@@
where IP represents the unique information pair generated through iterative refinement.
To further improve the quality and accuracy of reasoning results, ISP2 is designed to be compatible with different CoT prompting methods. Whether the task requires simple step-by-step reasoning or more sophisticated in-context learning, ISP2 dynamically integrates the appropriate prompting strategy to optimize performance. We illustrate an example of ISP2 reasoning in the figure above.
IV Experiments
@@ IV-A Experimental Setup
We evaluate $\text{ISP}^{2}$ on six datasets with diverse input formats. Extensive experiments are conducted across these datasets to demonstrate the universality of $\text{ISP}^{2}$ prompts. The Table I provides relevant information about the datasets used in our experiment, detailing the data source, task type, answer type, number of prompt samples, and total test samples for each dataset.
  • Mathematical Reasoning: The following four mathematical reasoning benchmarks are widely recognized and considered in the field: AddSub [35], which includes math word problems on addition and subtraction tailored for third to fifth graders; SVAMP [36], known for its math word problems with diverse structures; AQuA [37], which focuses on algebraic word problems; GSM8K [38], a published benchmark that features grade-school math problems.
  • Commonsense Reasoning: StrategyQA (SQA; [39]) and CommonsenseQA (CSQA; [40]) are utilized for commonsense tasks. CSQA comprises questions that require a variety of commonsense knowledge, while StrategyQA includes questions that necessitate multi-step reasoning.
TABLE I: Overview of datasets utilized in experiments
@@
GSM8K [38] Arithmetic Number 4 1,319
AddSub [35] Arithmetic Number 4 395
SVAMP [36] Arithmetic Number 4 1,000
AQuA [37] Arithmetic Multi-choice 4 254
StrategyQA [39] Commonsense True/False 4 2,290
CommonsenseQA [40] Commonsense Multi-choice 4 1,221
Note: “Number” represents the number of sampled datasets, and “Example” is the number of prompt examples in the same dataset.
@@
To effectively evaluate our method, we assess $\text{ISP}^{2}$ performance on three baseline prompting methods: Chain-of-Thoughts (CoT) [8], Complex CoT [12], Program-of-Thoughts (PoT) [41] and Self Consistency [10]. CoT predicts answers by generating explanations and steps, allowing the model to solve problems through explicit reasoning processes, which makes the decision-making more transparent. Complex CoT utilizes a complexity-based strategy, which breaks down complex problems into smaller parts, thereby improving reliability in complex scenarios. PoT prompts LLMs to generate executable code for problem decomposition and delegates computational steps to external interpreters, thereby disentangling reasoning from numerical processing to enhance accuracy and efficiency in complex tasks. Self Consistency generates multiple chains of thought and selects the highest-voted result as the final outcome through voting. Additionally, all experiments are conducted in a few-shot setting without training or fine-tuning the LLMs.
@@
We primarily use GPT-3.5, GPT-4o-mini, LLaMA2-7B, LLaMA2-13B, LLaMA3-8B, and DeepSeek-Distill-Qwen-7B for testing. Our decoding strategy employs greedy decoding with the temperature fixed at 0, ensuring deterministic outputs to eliminate stochasticity in comparisons between baseline methods and ISP2-enhanced variants. It minimizes the risk of hallucinations (e.g., generating inconsistent or nonsensical reasoning steps) while aligning with our experimental goal of isolating the intrinsic impact of ISP2. By fixing the temperature at 0, we ensure comparability across all methods and avoid conflating the effects of ISP2 with those of decoding stochasticity.
For the few-shot setting, we use four exemplars for most datasets, adjusting the number based on task complexity. We integrate ISP2 into baseline methods to evaluate its impact, denoted as ISP2-CoT, ISP2-CoT@5, ISP2-ComCoT, ISP2-ComCoT@5, ISP2-PoT, and ISP2-PoT@5. Here, ISP2-CoT, ISP2-ComCoT, and ISP2-PoT represent ISP2 combined with Chain-of-Thought (CoT), Complex CoT, and Program-of-Thoughts (PoT), respectively. The “@5” notation indicates the use of Self-Consistency (SC) by aggregating five reasoning chains via majority voting. Despite temperature=0, SC@5 introduces diversity through parallel sampling of distinct reasoning paths (enabled by varying the random seed in API calls), compensating for the reduced per-sample stochasticity. Additionally, we design task-specific answer format instructions in prompts to regulate the structure of final outputs (e.g., “Answer: [X]”) for precise answer extraction. Notably, the ISP2-PoT method specifically selects LLMs with strong code-generation capabilities (e.g., GPT and DeepSeek). By decoupling program derivation from code execution, it avoids the problem of solution failure caused by the model’s insufficient code implementation ability.
IV-B Main Results
TABLE II: The comparison results of existing SOTA plug-and-play methods on GPT-3.5.
GPT-3.5 turbo Prompt AddSub SVAMP GSM8K AQuA CSQA SQA Average
CoT [8] 81.2 81.2 76.2 58.7 75.5 69.7 73.75
PHP-CoT [42] 86.1 83.1 84.6 65.4 76.2 69.2 77.43
RE2-CoT [14] 82.7 84.9 81.2 63.3 77.9 67.1 76.60
ERA-CoT [43] 83.8 82.2 80.2 56.9 83.2 71.4 76.28
@@ IV-B Main Results
Note: We use accuracy as the evaluation metric.The best result is highlighted in bold, and the second best is underlined.
The main results are presented in Table III. Compared with the SOTA method that also functions as a plug-in, $\text{ISP}^{2}$ demonstrates considerable advantages. Significant improvements have been achieved on benchmarks such for GPT-4o-mini, GPT-3.5, Deepseek-Qwen-7B, LLaMA3-8B, LLaMA2-13B, and LLaMA2-7B. The average improvement rates were 1.1% for GPT-4o-mini, 7.1% for GPT-3.5, 4.5% for DeepSeek-Qwen-7B, 5.2% for LLaMA3-8B, 8.1% for LLaMA2-13B, and 12.4% for LLaMA2-7B. These results indicate that by processing $\text{ISP}^{2}$, LLM can better understand the essence of the problems and enhance its performance.
TABLE III: Result on Math Word Problem
@@
GPT-4o-mini CoT 96.1 93.4 92.3 81.4 90.80
$\text{ISP}^{2}$-CoT 97.2 93.9 92.4 83.7 91.80
+1.1 +0.5 +0.1 +2.3 +1.00
PoT 92.1 91.4 91.4 81.1 89.00
$\text{ISP}^{2}$-PoT 93.7 92.1 92.1 81.7 89.90
+1.6 +0.7 +0.7 +0.6 +0.90
ComCoT 96.7 93.8 91.8 82.3 91.15
$\text{ISP}^{2}$-ComCoT 97.2 94.1 92.3 83.7 91.83
+0.5 +0.3 +0.5 +1.4 +0.68
CoT@5 96.3 93.5 93.3 83.1 91.55
$\text{ISP}^{2}$-CoT@5 98.5 93.9 94.3 84.2 92.73
+2.2 +0.4 +1.0 +0.9 +1.18
PoT@5 95.6 94.1 92.2 81.7 90.90
$\text{ISP}^{2}$-PoT@5 97.0 94.3 92.8 82.7 91.70
+1.4 +0.2 +0.6 +1.0 +0.80
ComCoT@5 97.1 93.9 92.6 83.0 91.65
$\text{ISP}^{2}$-ComCoT@5 97.9 94.2 92.8 86.9 92.95
+0.8 +0.3 +0.2 +3.9 +1.30
GPT-3.5 Turbo CoT 81.2 81.2 76.2 58.7 74.33
$\text{ISP}^{2}$-CoT 88.6 88.7 83.4 64.4 81.28
+7.4 +7.5 +7.2 +5.7 +6.95
PoT 84.4 85.2 81.8 58.1 77.38
$\text{ISP}^{2}$-PoT 90.1 87.7 83.2 66.3 81.83
+5.7 +2.5 +1.4 +8.2 +4.45
ComCoT 82.7 80.1 79.3 57.8 74.98
@@
DeepSeek-
Distill-
Qwen-7B
CoT 89.1 86.9 75.3 47.2 74.63
$\text{ISP}^{2}$-CoT 90.5 91.4 82.8 48.9 78.40
+1.4 +4.5 +7.5 +1.7 +3.77
PoT 83.5 77.4 70.5 30.2 65.40
$\text{ISP}^{2}$-PoT 88.7 77.9 78.2 32.1 69.23
+5.2 +0.5 +7.7 +1.9 +3.83
ComCoT 88.4 93.4 78.2 62.6 80.65
$\text{ISP}^{2}$-ComCoT 91.5 94.3 86.3 65.8 84.48
+3.1 +0.9 +8.1 +3.2 +3.83
CoT@5 90.6 87.8 77.2 53.7 77.33
$\text{ISP}^{2}$-CoT@5 92.2 91.8 83.1 56.5 80.90
+1.6 +4.0 +5.9 +2.8 +3.57
PoT@5 84.4 80.0 71.8 32.3 67.13
$\text{ISP}^{2}$-PoT@5 92.5 82.5 79.8 33.3 69.78
+8.1 +2.5 +8.0 +1.0 +2.65
ComCoT@5 90.6 95.1 79.8 67.3 83.20
$\text{ISP}^{2}$-ComCoT@5 93.4 95.6 87.2 67.9 86.02
+2.8 +0.4 +9.9 +0.6 +2.82
LLaMA3-8B CoT 75.3 66.8 45.9 22.7 52.68
$\text{ISP}^{2}$-CoT 75.6 67.9 49.1 23.8 54.10
+0.3 +1.1 +3.2 +1.1 +1.42
ComCoT 75.6 66.2 49.7 25.0 54.13
$\text{ISP}^{2}$-ComCoT 76.5 73.0 50.6 27.8 56.98
+0.9 +6.8 +0.9 +2.8 +2.85
CoT@5 76.7 69.4 47.6 24.1 54.45
$\text{ISP}^{2}$-CoT@5 77.8 69.6 49.3 26.4 55.78
+1.1 +0.2 +1.7 +2.3 +1.33
ComCoT@5 77.1 66.4 50.3 26.4 55.05
$\text{ISP}^{2}$-ComCoT@5 77.5 74.8 51.7 27.9 57.98
+0.4 +8.4 +1.4 +1.5 +2.93
LLaMA2-13B CoT 38.1 36.3 16.7 16.6 26.93
$\text{ISP}^{2}$-CoT 60.0 39.2 19.3 20.4 34.73
+21.9 +2.9 +2.6 +3.8 +7.80
ComCoT 32.2 33.9 15.5 19.7 25.33
@@
Note: The @5 notation indicates usage of Self-Consistency with five reasoning chains for majority voting. ”ComCoT” stands for Complex CoT. ”PoT” stands for Program-of-Thoughts.
@@
Table III reports performance on the Math Word Problem (MWP) task, where our method achieves significant performance improvements across various mathematical subdomains. Our method not only achieves performance improvements in standard reasoning approaches, such as CoT, but also enables LLMs to enhance their performance through code generation. Notably, the enhancement is particularly significant when combined with Self Consistency, and the robustness and accuracy of $\text{ISP}^{2}$ on different models are evident. The AddSub dataset focuses on basic mathematical operations, and for models like GPT and Llama, they already possess sufficient capabilities to solve most problems. However, problems that are not successfully solved usually contain misleading information, and $\text{ISP}^{2}$ can avoid the influence of misleading information by filtering out unnecessary data in adaptive extraction, thus enabling more successful problem solving on LLMs. SVAMP and GSM8K are more advanced MWP datasets that focus on generalization capabilities in arithmetic and more complex algebra and geometry problems. Interestingly, we observe that the improvements of $\text{ISP}^{2}$ on SVAMP and GSM8K are not less than those on AddSub. $\text{ISP}^{2}$ can correct wrong directions through adaptive extraction and prevent misleading information from contaminating the summary process. Unlike the first three datasets, AQuA is an algebraic multiple-choice dataset, which not only requires the model to solve the problem but also to select the correct answer from multiple options. The format demands higher judgment capabilities from the model. $\text{ISP}^{2}$ facilitates further computations by allowing algebraic information pairs to be stored in formulaic forms. $\text{ISP}^{2}$ demonstrates good performance on mathematical datasets, enhancing the effectiveness of solutions by revealing key data essential for problem resolution, thus improving the performance of both ComplexCoT and Self Consistency.
@@
Notably, as detailed in Table II, compared to the current SOTA plug-and-play methods in mathematics, our approach performs well, reaching new best levels in AddSub and SVAMP, and second-best levels in GSM8K and AQuA. The main reason is that $\text{ISP}^{2}$ continuously improves solutions by selectively gathering previous descriptions of problem space, allowing it to accurately resolve issues. Instead of relying on multiple rounds of interactive reasoning, repetitive reading, or relationship extraction to obtain crucial information, solving many complex mathematical problems emphasizes the thought process and information filtering. By focusing on these aspects, misleading information is excluded, and effective problem-solving strategies are developed. It enables $\text{ISP}^{2}$ to achieve significant improvements in mathematical reasoning.
@@
Figure 4: Result on Commonsense Problem.
As shown in Figure 4, for the StrategyQA and CommonsenseQA datasets, $\text{ISP}^{2}$ has increased performance by $4.9\%$ and $2.8\%$, respectively. Additionally, Self Consistency maintains an orthogonal relationship with $\text{ISP}^{2}$, both enhancing the performance of CoT, enabling it to better understand common sense and make more accurate choices in commonsense reasoning. There is a significant performance improvement in the CSQA dataset, where many hidden pieces of information exist within the problems, and relying solely on explicit knowledge is insufficient for accurate responses. Notably, in StrategyQA with the LLaMA2 model, $\text{ISP}^{2}$ did not demonstrate significant improvement. Smaller parameter LLMs have a tendency to include irrelevant text when forming information pairs. Consequently, this inclusion leads to the formation of continuous erroneous reasoning chains. Such errors can impede the ability of $\text{ISP}^{2}$ to effectively enhance performance. However, most experiments still show that $\text{ISP}^{2}$ outperforms most base prompts in commonsense datasets, demonstrating strong capability of $\text{ISP}^{2}$ to guide CoT. When compared with existing SOTA methods, it performs comparably to ERA in SQA, which excels at controlling the reasoning process using entity relations, and also maintains second-best performance in CSQA.
IV-C Ablation Studies
TABLE IV: Ablation study on $\text{ISP}^{2}$ aiding CoT reasoning.